1) Determine the wavelength of light emitted when an electron in a hydrogen atom makes a transition from an orbital in n = 6 to an orbital in n = 5.

2) Determine the wavelength of the light absorbed when an electron in a hydrogen atom makes a transition from an orbital in which n = 2 to an orbital in which n =3
Answer #1

Answer (1) (2)

By using Rydberg equation,

\Delta E = R_{H}[(1/n_{i}^{2})-(1/n_{f}^{2})]

where RH = Rydberg Constant 2.18\times 10-18 J

nf = final energy level

ni = initial energy level

Therefore,

1) transition from n=6 to n=5;

nf = 5

ni = 6

\Delta E = 2.18 *\times10-18 J [(1/62) - (1/52)] = -2.66 * 10-20 J (-ve Sign represents energy released)

we know,

\Delta E = hc/\lambda  

where h= 6.626 \times10-34 Js and c= 3\times108 m/s

so, \lambda = (6.626\times10-34 Js \times 3\times108 m/s) / (2.66 * 10-20 J) = 7.47 \times 10-6 m = 7472nm

2) transition from n=2 to n=3;

nf = 3

ni = 2

\Delta E = 2.18 *\times10-18 J [(1/22) - (1/32)] = 3.03 * 10-19 J

\lambda = (6.626\times10-34 Js \times 3\times108 m/s) / (3.03 \times 10-19 J) = 6.56 \times 10-7 m = 656nm